EQUILIBRIUM STRUCTURES FOR THE CIS and TRANS ISOMERS OF 1,2-DIFLUOROETHYLENE AND THE CIS,TRANS ISOMER OF 1,4-DIFLUOROBUTADIENE

\underline{NORMAN C. CRAIG}, Department of Chemistry, Oberlin College, Oberlin, OH 44074; PE-TER GRONER, Department of Chemistry, University of Missouri-Kansas City, 5100 Rockhill Rd, Kansas City, MO 64110-2499; DONALD C. McKEAN, School of Chemistry, University of Edinburgh, West Mains Road, Edinburgh EH9 3JJ, U.K; MICHAEL J. TUBERGEN, Department of Chemistry, Kent State University, Kent, OH 44242.

Microwave spectra for seven isotopomers of cis-1,2-difluoroethylene have been reinvestigated and extended with the pulsedmolecular-beam Fourier-transform method. Rotational Hamiltonians that include a full set of quartic centrifugal distortion constants have been fit for each isotopomer. Ground state rotational constants for the trans isomer were available from an investigation of the high-resolution infrared spectra of several isotopomers. Equilibrium rotational constants have been extracted for the cis and trans isomers from the experimental ground state constants and the vibration-rotation constants computed at the MP2/6-31G(d) level. Equilibrium structures have been fit to the equilibrium principal moments of inertia for both isotopomers, and isomeric differences in the bond lengths and bond angles are interpreted in terms of the strong influence of fluorine substitution. For cis,trans-1,4-difluorobutadiene a similar derivation of the equilibrium structure has been achieved through the combination of published microwave data and quantum chemical calculations. This substance provides an opportunity to examine the influence of cis and trans fluorine substitution within a single molecule.